OpenStudy (anonymous):

Medal? Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3.

OpenStudy (anonymous):

y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3

OpenStudy (anonymous):

@aliboo997 Can you help?

OpenStudy (anonymous):

Only one of your options is in point slope form.

OpenStudy (anonymous):

Point-Slope Form: y - y1 = m(x - x1)

OpenStudy (anonymous):

So would it be B?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

Thank you